OpenStudy (anonymous):

Please Help I Give Medals XD A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi

ganeshie8 (ganeshie8):

2x + 3y = 1,200

ganeshie8 (ganeshie8):

start by calculating \(slope\) of this line

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

to find the \(slope\) , you need to convert it to slope-intercept form first :- y = mx + b

ganeshie8 (ganeshie8):

m = slope b = y-intercept

ganeshie8 (ganeshie8):

see if u can rearrange the given equation to that form

OpenStudy (anonymous):

X and Y are unknown though : / sorry I really suck at math

ganeshie8 (ganeshie8):

no, u dont... its easy let me show u. we're just rearranging :- \( 2x + 3y = 1200\) subtract 2x both sides \(3y = -2x + 1200\) divide 3 both sides \(y = -\frac{2}{3}x + 400\)

ganeshie8 (ganeshie8):

compare it wid the slope intercept form :- \(y = mx + b\)

ganeshie8 (ganeshie8):

\(m = ?\) \(b = ?\)

OpenStudy (anonymous):

m is the slope so...

ganeshie8 (ganeshie8):

yes so...

OpenStudy (anonymous):

Dude theres no hope for me with math, I'm sorry for wasting your time but I'm just to stupid to understand it. So I'm gonna give you a medal and I won't bother you again...

ganeshie8 (ganeshie8):

yo dont give up so easily+cheaply lol

ganeshie8 (ganeshie8):

these are easy if u persist a bit longer time

ganeshie8 (ganeshie8):

earlier, we rarranged the given equation to slope-intercept form :- \(\large y = -\frac{2}{3}x + 400 \) comparing it wid the equation :- \(\large y = mx + b\) so, \(\large m = -\frac{2}{3}\) \(\large b = 400\)

ganeshie8 (ganeshie8):

I have just compared both the equations, now I have \(m\) and \(b\), so I can start graphing the line now.

ganeshie8 (ganeshie8):

see if things make more or less sense so far ?

OpenStudy (anonymous):

more actually :)

ganeshie8 (ganeshie8):

good :) lets graph the equation. There is a simple way to graph the line, when u knw \(slope\) and \(y-intercept\)

ganeshie8 (ganeshie8):

here are the steps :- step1 : plot the \(y-intercept\)

ganeshie8 (ganeshie8):